# ED4: Explicit Data-level Debiasing for Deepfake Detection

### v1

@@ Introduction
As shown in Fig. 2, we propose the Adversarial Spatial Consistency Module (AdvSCM) to address spatial bias in the extracted feature. The adversarial consistency strategy refers to the generator (enlarging spatial inconsistency) vs. the backbone extractor (reducing feature inconsistency), which closely resembles the generator vs. discriminator in the generative adversarial network. By achieving this, the backbone can learn to extract features that are spatial-agnostic, thus avoiding spatial bias. In AdvSCM, we first introduce spatial shuffle (Chen et al. 2019) as the base operation, which does not undermine the texture and fine-grained information but only produces spatial inconsistency. Then, given an input image $\mathbf{I}$, two shuffled versions of $\mathbf{I}$ are generated by the random generator $\mathcal{G}_{r}(\cdot)$ and the adversarial generator $\mathcal{G}_{a}(\cdot;\theta_{a})$, respectively, where $\theta_{a}$ is the learnable parameter in $\mathcal{G}$. Without learnable parameters, $\mathcal{G}_{r}$ relocate blocks and reassemble them to obtain the shuffled image $\mathbf{I}_{s1}=\mathcal{G}_{r}(\mathbf{I})$ with random granularity and permutation. In contrast, $\mathcal{G}_{a}$ is a neural network that inputs the concatenation of $\mathbf{I}$ and $\mathbf{I}_{s1}$, and outputs a probability metric $\mathbf{m}=\mathcal{G}_{a}((\mathbf{I},\mathbf{I}_{s1});\theta_{a})\in\mathbb{R}^{N\times N}$, where $N$ is the number of divided patches for shuffling. $\mathbf{m}$ is then transformed via Hungarian algorithm (Kuhn 1955) to $\mathbf{\hat{m}}\in\{0,1\}^{N\times N}$, which can be treated as the shuffle permutation for generating $\mathbf{I}_{s2}$. More details can be found in Supplementary Material. Both $\mathbf{I}_{s1}$ and $\mathbf{I}_{s2}$ are then input into the backbone extractor $\mathcal{E}(\cdot;\theta_{e})$ to obtain their respective features $\mathbf{F}_{s1},\mathbf{F}_{s2}$. The optimization target of $\theta_{e}$ is minimizing the distance between $\mathbf{F}_{s1}$ and $\mathbf{F}_{s2}$, which can be formulated as:
@@
where $\theta_{e}$ should be frozen. Notably, the gradient flow for optimizing $\theta_{a}$ is broken by non-differentiable operations like sampling and the Hungarian algorithm. Therefore, we approximate the gradient calculation via Reinforce algorithm (Williams 1992):
@@
By introducing AdvSCM, the adversarial generator learns to predict shuffle configuration that maximizes the spatial inconsistency between generated shuffled images, thus allowing the backbone extractor to more robustly learn spatial-agnostic features via consistency constraints.
Detection Loss
@@
Experimental Results
Experimental Setting
@@
Datasets Given that the generalization issue is the major challenge for research, we apply an abundant number of advanced and widely used deepfake datasets in our experiments. FaceForensics++ (FF++) (Rossler et al. 2019) is constructed by four forgery methods including Deepfakes (DF) (DeepFakes. 2021), Face2Face (F2F) (Thies et al. 2016), FaceSwap (FS) (FaceSwap. 2021), and NeuralTextures (NT) (Thies, Zollhöfer, and Nießner 2019). Meanwhile, it includes three different compression quality levels, that is, RAW, High-Quality (HQ), and Low-Quality (LQ). We employ FF++ (HQ) as the training dataset for all experiments in our paper. For cross-dataset evaluations, we introduce Celeb-DF-v1 (CDFv1) (Li et al. 2020b), Celeb-DF-v2 (CDFv2) (Li et al. 2020b), DeepFakeDetection (DFD) (DFD. 2021), and DeepFake Detection Challenge (DFDC) (detection challenge. 2021).
@@
Implementation Details For preprocessing and training, we strictly follow the official code and settings provided by DeepFakeBench (Yan et al. 2023b) to ensure fair comparison. For ClockMix, the number of mixing images is selected from $\{1,2,3,4\}$, and $\rho$ is randomly selected from 45 to 315. Our method takes Xception (Chollet 2017) as the backbone for both adversarial generator and extractor and they are initialized by the parameters pre-trained on ImageNet (Krizhevsky, Sutskever, and Hinton 2012). The Adam optimizer is used with a learning rate of 0.0002, epoch of 10, input size of 256 $\times$ 256, and batch size of 32. Frame-level Area Under Curve (AUC) and Equal Error Rate (EER) (Yan et al. 2023b) are applied as the evaluation metrics of experimental results. All experiments are conducted on one NVIDIA Tesla V100 GPUs.
Table 1: Cross-dataset evaluations (AUC) from FF++ to CDFv1, CDFv2, DFD and DFDC. C-Avg. denotes the average value of cross-dataset results. The best results are highlighted in bold.
Method Venue FF++ CDFv1 CDFv2 DFD DFDC C-Avg.
Xception (Chollet 2017) CVPR’17 0.9637 0.7794 0.7365 0.8163 0.7077 0.7600
FWA (Li and Lyu 2018) CVPRW’18 0.8765 0.7897 0.6680 0.7403 0.6132 0.7028
Meso4 (Afchar et al. 2018) WIFS’18 0.6077 0.7358 0.6091 0.5481 0.5560 0.6123
Capsule (Nguyen, Yamagishi, and Echizen 2019) ICASSP’19 0.8421 0.7909 0.7472 0.6841 0.6465 0.7172
EfficientNetB4 (Tan and Le 2019) ICML’19 0.9567 0.7909 0.7487 0.8148 0.6955 0.7625
CNN-Aug (Wang et al. 2020) CVPR’20 0.8493 0.7420 0.7027 0.6464 0.6361 0.6818
X-ray (Li et al. 2020a) CVPR’20 0.9592 0.7093 0.6786 0.7655 0.6326 0.6965
FFD (Dang et al. 2020) CVPR’20 0.9624 0.7840 0.7435 0.8024 0.7029 0.7582
F3Net (Qian et al. 2020) ECCV’20 0.9635 0.7769 0.7352 0.7975 0.7021 0.7530
SPSL (Liu et al. 2021) CVPR’21 0.9610 0.8150 0.7650 0.8122 0.7040 0.7741
SRM (Luo et al. 2021) CVPR’21 0.9576 0.7926 0.7552 0.8120 0.6995 0.7648
CORE (Ni et al. 2022) CVPRW’22 0.9638 0.7798 0.7428 0.8018 0.7049 0.7573
Recce (Cao et al. 2022) CVPR’22 0.9621 0.7677 0.7319 0.8119 0.7133 0.7562
UIA-VIT (Zhuang et al. 2022) ECCV’22 0.9655 0.7925 0.8002 0.8279 0.7038 0.7811
IID (Huang et al. 2023) CVPR’23 0.9743 0.7578 0.7687 0.7935 0.6951 0.7538
UCF (Yan et al. 2023a) ICCV’23 0.9705 0.7793 0.7527 0.8074 0.7191 0.7646
Liang et al.(Liang, Shi, and Deng 2022) ECCV’22 - 0.706 - 0.829 0.700 -
SLADD (Chen et al. 2022) CVPR’22 0.9691 0.8015 0.7403 0.8089 0.7170 0.7669
SBI (Shiohara and Yamasaki 2022) CVPR’22 0.9747 0.8371 0.7911 0.8112 0.7275 0.7917
LSDA (Yan et al. 2024) CVPR’24 - 0.867 0.830 0.880 0.736 0.828
Ours - 0.9806 0.8871 0.8394 0.8765 0.7427 0.8364
Cross-dataset Evaluation
In Tab. 1, we provide extensive comparison results with existing state-of-the-art (SOTA) deepfake detectors based on DeepFakeBench (Yan et al. 2023b). Specifically, all methods are trained on FF++ (HQ) and tested on other datasets. The methods in the upper part of the table are within the benchmark with the exact same experimental setting as our method. Hence, we directly copy their results from DeepfakeBench. Notably, more SOTA methods are related to our method but are not included in the DeepFakeBench. Considering their implementation details are various (such as SBI (Shiohara and Yamasaki 2022) is trained with FF++ RAW, size of 384, and tested with video-level AUC), we carefully reproduce SLADD (Chen et al. 2022) and SBI following their official code within the DeepFakeBench framework, and the results of Liang et al. (Liang, Shi, and Deng 2022) are reproduced by (Yan et al. 2023a). LSDA (Yan et al. 2024) is the new SOTA based on DeepFakeBench, therefore we directly copy the results from their paper.
@@
It can be observed that our method consistently outperforms the earlier deepfake detectors based on data synthesis techniques (i.e., FWA, X-ray, SLADD, and SBI) across all evaluated datasets. Meanwhile, compared with the implicit disentanglement methods (i.e., Liang et al.and UCF), our method also exhibits superior effectiveness. More cross-dataset results for our method can be found in Supplementary Material. These results demonstrate the superiority of the proposed explicit data-level debiasing approach over previous SOTA methods.
Ablation Study
@@
Table 2: The impact of ClockMix (CM) and Adversarial Spatial Consistency Module (AdvSCM).
@@
In Tab. 2, we assess the effectiveness of each proposed component. The detector trained without all proposed components is denoted by Baseline. The results on all evaluated datasets and metrics demonstrate that each component essentially contributes to the effectiveness of our method.
Table 3: Comparison of different mixing and label-assignment strategies. HL indicates assigning hard labels with Eq. (5), and SL denotes calculating soft labels with Eq. (6). MRF denotes assigning the mixture of real images with the fake label.
@@
To demonstrate the superiority of the proposed ClockMix, we replace ClockMix with traditional data mixing augmentations, that is, Mosaic (Bochkovskiy, Wang, and Liao 2020), Mixup (Zhang et al. 2017) and CutMix (Yun et al. 2019). We also investigate the influences of label-assigning strategies, including Hard Label (HL), Soft Label (SL), and assigning the Mixture of Real images with the Fake label (MRF). In Tab. 3, we first observe that MRF fundamentally destroys the detector’s ability to justify forgery clues. This can be attributed to all images with ClockMix being assigned with Fake labels, hence the detector mistakenly takes the ClockMix splicing patterns as the indicator of forgery. The proposed ClockMix-HL significantly outperforms all ablation variants in performance. Image with Mosaic cannot be applied with HL considering the intact information is preserved for each mixed face. Meanwhile, it can be observed that SL severely undermines the effectiveness of CutMix and ClockMix by categorizing the Real+Fake images under an uncertain label, which reduces the confidence in determining fake images with minor forgery artifacts. Since applying Mixup renders the details of forgery artifacts, both HL and SL versions of Mixup compromise the effectiveness of the deepfake detector.
Table 4: Effectiveness of the proposed adversarial strategy compared with random strategies.
@@
To validate the effectiveness of the proposed adversarial consistency strategy, we design two ablation variants: 1) Random Shuffle (RS), where we only deploy random shuffle to the image, and take the shuffled image as the training input. 2) Random Shuffle Consistency (RS+C), where we generate two random shuffle versions of an image, and apply Eq. (7) to encourage their consistency without using Eq. (9) for the adversarial strategy. Tab. 4 shows that RS+C exhibits slightly superior effectiveness on the DFD dataset. However, it cannot stably perform best in different cross-dataset settings. In contrast, deploying AdvSCM enhances the overall cross-dataset generalization ability of deepfake detection.

### v2

@@ IV-C Adversarial Spatial Consistency Module
As shown in Fig. 3, we propose the Adversarial Spatial Consistency Module (AdvSCM) to address spatial bias in the extracted feature. The adversarial consistency strategy refers to the generator (enlarging spatial inconsistency) vs. the backbone extractor (reducing feature inconsistency), which closely resembles the generator vs. discriminator in the generative adversarial network. By achieving this, the backbone can learn to extract features that are spatial-agnostic, thus avoiding spatial bias. In AdvSCM, we first introduce spatial shuffle shuffle as the base operation, which does not undermine the texture and fine-grained information but only produces spatial inconsistency. Then, given an input image $\mathbf{I}$, two shuffled versions of $\mathbf{I}$ are generated by the random generator $\mathcal{G}_{r}(\cdot)$ and the adversarial generator $\mathcal{G}_{a}(\cdot;\theta_{a})$, respectively, where $\theta_{a}$ is the learnable parameter in $\mathcal{G}$. Without learnable parameters, $\mathcal{G}_{r}$ relocate blocks and reassemble them to obtain the shuffled image $\mathbf{I}_{s1}=\mathcal{G}_{r}(\mathbf{I})$ with random granularity and permutation. In contrast, $\mathcal{G}_{a}$ is a neural network that inputs the concatenation of $\mathbf{I}$ and $\mathbf{I}_{s1}$, and outputs a probability metric $\mathbf{m}=\mathcal{G}_{a}((\mathbf{I},\mathbf{I}_{s1});\theta_{a})\in\mathbb{R}^{N\times N}$. Namely, given a granularity $g$, the number of divided patches is $N=g^{2}$. To be more specific, $g$ represents the number of divisions along both the horizontal and vertical axes of an image, such that the image is then divided into $N=g^{2}$ patches. Hence, the shape of the output possibility metric $\mathbf{m}$ is $N\times N$, where the first $N$ refers to the number of patches in the shuffled image, the second one refers to the possible shuffled location of patches. Such an effect can be easily achieved by deploying a linear layer that outputs $N\times N$ after the network backbone and then reshaping the output. In implementation, the feature extraction backbone of $G_{a}$ remains shared across all different $N$. Only the number of channels in the final linear layer is adjusted to accommodate the current $N$ value. However, directly using $\mathbf{m}$ to guide the shuffling of images could be problematic. Specifically, each region should be assigned to a unique patch, that is, the predicted region index for each patch must be distinct. In fact, this mutual exclusivity of indices aligns with solving polynomials, except we are maximizing polynomials instead of minimizing them, which could be easily transformed by adding a minus sign to $\mathbf{m}$. Therefore, the problem can be formulated as:
where $x_{i,j}\in\{0,1\}$ and
Figure 6: Visualized example of transforming generator output to shuffle permutation via linear programming.
Both $\mathbf{I}_{s1}$ and $\mathbf{I}_{s2}$ are then input into the backbone extractor $\mathcal{E}(\cdot;\theta_{e})$ to obtain their respective features $\mathbf{F}_{s1},\mathbf{F}_{s2}$. The optimization target of $\theta_{e}$ is minimizing the distance between $\mathbf{F}_{s1}$ and $\mathbf{F}_{s2}$, which can be formulated as:
@@
where $\theta_{e}$ should be frozen. Notably, the gradient flow for optimizing $\theta_{a}$ is broken by non-differentiable operations like sampling and the Hungarian algorithm. Therefore, we approximate the gradient calculation via the Reinforce algorithm reinforce :
@@
By introducing AdvSCM, the adversarial generator learns to predict a shuffle configuration that maximizes the spatial inconsistency between generated shuffled images, thus allowing the backbone extractor to more robustly learn spatial-agnostic features via consistency constraints. The superiority of the proposed AdvSCM could be summarised as follows:
  • First Explicit Spatial-Bias Mitigation via Shuffle Consistency and Adversarial Adaptation. AdvSCM is the first method explicitly designed to counteract the detector’s central-region bias: it employs a shuffle-consistency strategy to disrupt over-reliance on structural forgery cues at the facial center and integrates an adversarial module to compel robust adaptation to manipulations at arbitrary spatial locations.
  • Multi-Granularity Feature Consistency Across Rearrangements. Unlike prior spatial-aware approaches that rely on implicit spatial transformations, relational constraints, or frame thumbnail rearrangements, AdvSCM enforces consistency at multiple feature scales between two distinct spatial permutations of the same image, thereby more effectively disentangling genuine content features from positional bias.
TABLE I: Cross-dataset evaluations (AUC) from FF++ to CDFv1, CDFv2, DFD and DFDC based on DeepfakeBench deepfakebench (Protocol 1). C-Avg. denotes the average value of cross-dataset results. The best results are highlighted in bold.
Method Venue FF++ FF++ CDFv1 Celeb-df CDFv2 Celeb-df DFD dfd DFDC dfdc C-Avg.
Xception xception CVPR’17 0.9637 0.7794 0.7365 0.8163 0.7077 0.7600
FWA FWA CVPRW’18 0.8765 0.7897 0.6680 0.7403 0.6132 0.7028
Meso4 mesonet WIFS’18 0.6077 0.7358 0.6091 0.5481 0.5560 0.6123
Capsule capsule ICASSP’19 0.8421 0.7909 0.7472 0.6841 0.6465 0.7172
EfficientNetB4 effnet ICML’19 0.9567 0.7909 0.7487 0.8148 0.6955 0.7625
CNN-Aug cnn-Aug CVPR’20 0.8493 0.7420 0.7027 0.6464 0.6361 0.6818
X-ray xray CVPR’20 0.9592 0.7093 0.6786 0.7655 0.6326 0.6965
FFD ffd CVPR’20 0.9624 0.7840 0.7435 0.8024 0.7029 0.7582
F3Net F3Net ECCV’20 0.9635 0.7769 0.7352 0.7975 0.7021 0.7530
SPSL spsl CVPR’21 0.9610 0.8150 0.7650 0.8122 0.7040 0.7741
SRM srm CVPR’21 0.9576 0.7926 0.7552 0.8120 0.6995 0.7648
CORE core CVPRW’22 0.9638 0.7798 0.7428 0.8018 0.7049 0.7573
Recce recce CVPR’22 0.9621 0.7677 0.7319 0.8119 0.7133 0.7562
SLADD adv CVPR’22 0.9691 0.8015 0.7403 0.8089 0.7170 0.7669
SBI SBI CVPR’22 0.9747 0.8371 0.7911 0.8112 0.7275 0.7917
UIA-VIT uia ECCV’22 0.9655 0.7925 0.8002 0.8279 0.7038 0.7811
IID huang2023implicit CVPR’23 0.9743 0.7578 0.7687 0.7935 0.6951 0.7538
UCF ucf ICCV’23 0.9705 0.7793 0.7527 0.8074 0.7191 0.7646
Liang et al.Disentangle ECCV’22 - 0.706 - 0.829 0.700 -
LSDA lsda CVPR’24 - 0.867 0.830 0.880 0.736 0.828
ED4 (Ours) - 0.9806 0.8871 0.8394 0.8765 0.7427 0.8364
TABLE II: Cross-dataset evaluations on more various and advanced datasets (Protocol 2). The metric is frame-level AUC and all models are trained on FF++ (HQ) FF++ .
Methods WDF FAVC DiffSwap E4S BlendFace UniFace DFL SDv21 Avg. #Top1
EfficientNet effnet 0.7275 0.8404 0.7956 0.6514 0.7813 0.7775 0.7347 0.5581 0.7333 0
Xception xception 0.7705 0.8681 0.8110 0.6554 0.7911 0.8282 0.7199 0.6145 0.7573 0
SPSL spsl 0.7023 0.7689 0.7784 0.5987 0.6936 0.6915 0.6548 0.5877 0.6845 0
UCF ucf 0.7738 0.8631 0.8511 0.6916 0.7850 0.7867 0.7433 0.5910 0.7607 1
DCL dcl 0.7401 0.6169 0.8245 0.7611 0.7992 0.7145 0.7394 0.6315 0.7284 1
TALL tall 0.7428 0.8031 0.7971 0.7030 0.8320 0.8310 0.6761 0.6501 0.7544 0
CDFA cdfa 0.7437 0.8861 0.5479 0.6256 0.6840 0.7576 0.6139 0.6070 0.6832 0
RAM ram 0.6991 0.7742 0.5991 0.5991 0.8679 0.8323 0.7530 0.5071 0.7084 2
ED4 (Ours) 0.7682 0.8879 0.8781 0.7201 0.8511 0.8280 0.7794 0.6704 0.7979 5
IV-D Detection Loss
@@
V Experimental Results
V-A Experimental Setting
Datasets
@@ V-A Experimental Setting
Given that the generalization issue is the major challenge for research, we apply an abundance of advanced and widely used deepfake datasets in our experiments. FaceForensics++ (FF++) FF++ is constructed by four forgery methods including Deepfakes (DF) df , Face2Face (F2F) f2f , FaceSwap (FS) fs , and NeuralTextures (NT) nt . Meanwhile, it includes three different compression quality levels, that is, RAW, High-Quality (HQ), and Low-Quality (LQ). We employ FF++ (HQ) as the training dataset for all experiments in our paper. For cross-dataset evaluations, we introduce both classical and advanced datasets, including Celeb-DF-v1 (CDFv1) Celeb-df , Celeb-DF-v2 (CDFv2) Celeb-df , DeepFakeDetection (DFD) dfd , DeepFake Detection Challenge (DFDC) dfdc , DF40 df40 , DiffusionFace diffusionface , and more. These datasets have covered all types of forgery categories, including Face-Swapping, Face-Reenactment, Face-Editing, and Entire Face Synthesis df40 .
Implementation Details
@@
For preprocessing and training, we strictly follow the official code and settings provided by DeepFakeBench deepfakebench to ensure fair comparison. DeepFakeBench is a comprehensive benchmark tailored for deepfake detection, which proposes a unified framework for pre-processing, training, and testing. They also reproduced the SoTA methods with consistent training hyper-parameters. To be specific, Dlib dlib is used for face extraction and alignment with a cropping margin of 15%. 32 frames are extracted from each video and all frames are cropped to the size of 256 $\times$ 256. For training, traditional data augmentations are deployed, including rotating, Gaussian noise, saturation adjusting, and quality adjusting. For ClockMix, the number of mixing images is randomly selected from $\{1,2,3,4\}$, and $\rho$ is randomly selected from 45 to 315. Our method takes Xception xception as the backbone for both adversarial generator and extractor and they are initialized by the parameters pre-trained on ImageNet imagenet . The Adam optimizer is used with a learning rate of 0.0002, an epoch of 10, an input size of 256 $\times$ 256, and a batch size of 32. Frame-level Area Under Curve (AUC) and Equal Error Rate (EER) deepfakebench are applied as the evaluation metrics of experimental results. All experiments are conducted on one NVIDIA Tesla V100 GPU.
V-B Cross-dataset Evaluation
Based on DeepFakeBench (Protocol 1)
@@ V-B Cross-dataset Evaluation
In Tab. I, we provide extensive comparison results with existing state-of-the-art (SOTA) deepfake detectors based on DeepFakeBench deepfakebench . Specifically, all methods are trained on FF++ (HQ) and tested on other datasets. The methods in the upper part of the table are within the benchmark with the exact same experimental setting as our method. Hence, we directly copy their results from DeepfakeBench. Notably, more SOTA methods are related to our method but are not included in the DeepFakeBench. Therefore, the results of Liang et al. Disentangle are reproduced by ucf . LSDA lsda is the new SOTA based on DeepFakeBench, hence we directly copy the results from their paper. It can be observed that our method consistently outperforms the earlier deepfake detectors based on data synthesis techniques (i.e., FWA, X-ray, SLADD, and SBI) across all evaluated datasets. Meanwhile, compared with the implicit disentanglement methods (i.e., Liang et al.and UCF), our method also exhibits superior effectiveness. These results demonstrate the superiority of the proposed explicit data-level debiasing approach over previous SOTA methods.
More Advanced Datasets (Protocol 2)
Here, we provide results of our method on seven more datasets, including WildDeepfake wilddeepfake (WDF), FakeAVCeleb fakeavceleb (FAVC), {E4S e4s , BlendFace blendface , UniFace uniface , DeepfaceLab deepfacelab (DFL)} from DF40 df40 , and {DiffSwap diffswap , SDv21 sd } from DiffusionFace diffusionface . Considering most baseline methods from DeepfakeBench deepfakebench have not been reproduced on these advanced datasets, we carefully reproduce EfficientNet effnet , Xception xception , SPSL spsl , UCF ucf , DCL dcl , TALL tall , CDFA cdfa , and RAM ram , and then evaluate them on these datasets for comparison. As shown in Tab. II, it can be observed that our methods can be generalized to various datasets with superior effectiveness.
V-C Ablation Study
@@
TABLE III: The impact of ClockMix (CM) and Adversarial Spatial Consistency Module (AdvSCM).
@@
In Tab. III, we assess the effectiveness of each proposed component. The detector trained without all proposed components is denoted by Baseline. The results on all evaluated datasets and metrics demonstrate that each component essentially contributes to the effectiveness of our method.
TABLE IV: Comparison of different mixing and label-assignment strategies. HL indicates assigning hard labels with Eq. (5), and SL denotes calculating soft labels with Eq. (6). MRF denotes assigning the mixture of real images with the fake label.
@@
To demonstrate the superiority of the proposed ClockMix, we replace ClockMix with traditional data mixing augmentations, that is, Mosaic yolov4 , Mixup mixup , and CutMix cutmix . We also investigate the influences of label-assigning strategies, including Hard Label (HL), Soft Label (SL), and assigning the Mixture of Real images with the Fake label (MRF). In Tab. IV, we first observe that MRF fundamentally destroys the detector’s ability to justify forgery clues. This can be attributed to all images with ClockMix being assigned with Fake labels, hence the detector mistakenly takes the ClockMix splicing patterns as the indicator of forgery. The proposed ClockMix-HL significantly outperforms all ablation variants in performance. Image with Mosaic cannot be applied with HL considering the intact information is preserved for each mixed face. Meanwhile, it can be observed that SL severely undermines the effectiveness of CutMix and ClockMix by categorizing the Real+Fake images under an uncertain label, which reduces the confidence in determining fake images with minor forgery artifacts. Since applying Mixup renders the details of forgery artifacts, both HL and SL versions of Mixup compromise the effectiveness of the deepfake detector.
TABLE V: Effectiveness of the proposed adversarial strategy compared with random strategies.
@@
TABLE VI: Comparison with more augmentation methods for spatial bias. Local represents the introduction of local forgery. Region denotes introducing region variations. Info-I denotes Informative pixel Integrity.
Method Local Region Info-I DFD CDFv2 DFDC Avg.
w/o AdvSCM (Baseline) 0.8200 0.7712 0.7296 0.7736
GridMask 0.8408 0.7838 0.7312 0.7853
AttrMask 0.8005 0.7139 0.6759 0.7301
RandomCrop 0.8415 0.8150 0.7235 0.7933
Shuffle 0.8765 0.8394 0.7427 0.8195
Figure 7: Robustness against unseen perturbations with multiple intensities. Block-wise, Blur, and Hue denote block-wise masking, median blur, and hue adjustment, respectively. The curves across multiple intensities consistently demonstrate that ED4 maintains higher performance than the baseline when confronted with unseen perturbations, thereby highlighting its improved practical applicability for deepfake detection.
Figure 8: Robustness against spatial deviations. “Variations" indicates the variation of prediction scores when applying different spatial transformations. The dotted lines represent the average variation. The results indicate the spatial bias issue in the previous baseline, and also demonstrate the effectiveness of ED4 in spatial bias removal.
@@
To validate the effectiveness of the proposed adversarial consistency strategy, we design two ablation variants: 1) Random Shuffle (RS), where we only deploy random shuffle to the image, and take the shuffled image as the training input. It could be treated as naively deploying Grid Shuffle as a data augmentation operation. 2) Random Shuffle Consistency (RS+C), where we generate two random shuffle versions of an image, and apply Eq. (9) to encourage their consistency without using Eq. (11) for the adversarial strategy. w/o AdvSCM could be treated as the baseline for this ablation. Tab. V shows that only applying RS contributes trivially to performance enhancement. RS+C could also improve the generalization performance of the detector, which further demonstrates that the shuffle consistency constraint could reduce the spatial bias in the extracted features. However, despite RS+C exhibiting slightly superior effectiveness on the DFD dataset, it cannot stably perform best in different cross-dataset settings. In contrast, deploying AdvSCM enhances the overall cross-dataset generalization ability of deepfake detection.
Comparison with More Augmentation Methods for Spatial Bias
As shown in Tab. VI, we introduce three approaches to replace Shuffle that may contribute to spatial bias removal: 1) GridMask gridmask , which is a widely-used augmentation that randomly masks a certain ratio of images. 2) AttrMask, which can mask out certain facial attributes guided by the facial landmarks. 3) RandomCrop, which alters the relative location of faces in the images. The w/o AdvSCM can be treated as the baseline for which no spatial transformation is deployed. The masking-based methods can break the structural forgery into local ones, but cannot improve the model sensitivity to artifacts at arbitrary regions. While RandomCrop can alter the forgery location to some extent, but still maintains the forgery structure. Meanwhile, since masking and cropping lead to the reduction of informative pixels, the performances achieved by these methods are not as pronounced as our S-Shuffle. Especially for AttrMask, despite masking certain facial attributes that may enhance the model sensitivity to local forgery, it significantly reduces the pixels that likely contain forgery information and thus leads to performance degradation. Moreover, this information reduction further leads to inherent inconsistency in extracted features, thereby rendering the employment of Consistency Loss inappropriate. In contrast, deploying Shuffle can uniformly address spatial bias while maintaining informative pixel integrity. Therefore, we use Shuffle as the basic spatial operation to construct the proposed adversarial spatial consistency module.
TABLE VII: Investigation for mixing images with different labels. Avg. denotes the Average performance in the cross-dataset evaluation.
Method DFD CDFv2 DFDC Avg.
Baseline 0.7994 0.7454 0.7089 0.7512
w/o Real+Real 0.7177 0.6959 0.6361 0.6832
w/o Real+Fake 0.8481 0.8045 0.7255 0.7927
w/o Fake+Fake 0.8560 0.7931 0.7247 0.7913
Ours 0.8765 0.8394 0.7427 0.8195
V-D The Effectiveness of Conducting ClockMix on Arbitrary Images
To investigate the influence of conducting ClockMix on arbitrary images, we perform an ablation study with three ablation variants, that is, w/o Real+Real. w/o Real+Fake, and w/o Fake+Fake. As shown in Tab. VII, w/o Real+Real will undermine the performance by indulging the detector to take the seaming patterns as fake. Such a result is similar to the “mixture of real images with the fake label (MRF)” that we present in the main paper except the degradation is slighter. Overall, the lack of any label combination in ClockMix can lead to considerable performance degradation, which indicates the superiority of mixing arbitrary images over existing label-specific and image-specific synthesis methods.
V-E Robustness Evaluation
Here, we evaluate the robustness of our method from two perspectives, that is, against unseen perturbations and against spatial deviations. Namely, perturbations may be easily applied to the images during data transportation, image compression, and bad light conditions of picture capturing. Spatial deviations are also common in real-world scenarios like unstable shooting positions or accidental data loss. All experiments for robustness are conducted on the FaceForensics++ dataset, and we take our backbone network Xception as the baseline.
Robustness Against Unseen Perturbations. As shown in Fig. 7, we introduce block-wise masking, median blur, and hue adjustment with multiple intensities for the robustness evaluation against unseen perturbations. To comprehensively investigate the robustness, we set the intensity of perturbations to gradually increase. Specifically, the perturbation intensities represent the masking ratios $\{0,0.05,0.10,0.15,0.20\}$ with fixed $8\times 8$ grids for Block-wise, the kernel sizes $\{0,3,4,5,6\}$ for Blur, and hue adjustment ratios $\{0,0.1,0.2,0.3,0.4\}$ for Hue. Obviously, our method exhibits superior robustness due to the improved reception of forgery clues.
Robustness Against Spatial Deviations. We provide a prediction stability study to investigate the robustness against spatial deviation. Specifically, we quantify the variation in prediction confidence for each image after applying the spatial deviations (i.e., Crop, Resize, and Shift) and subsequently depict these variations as a histogram in Fig. 8. Mathematically, the deviation value at each image can be written as $|\mathcal{N}(\mathbf{I})-\mathcal{N}(\mathcal{T}_{s}(\mathbf{I}))|$, where $\mathcal{N}(\cdot)$ is the detection network, $\mathbf{I}$ is the original image, and $\mathcal{T}_{s}$ is one of the spatial transforming operations. A larger variation indicates a worse detection stability. For the baseline detector, a considerable number of detection results suffer from spatial deviations. Especially for Crop, over one-third of the variations exceed 0.5, indicating a notable number of detection results are reversed to the opposite after spatial deviation. In contrast, with the help of AdvSCM, our detector can maintain strong robustness against different spatial deviations.